# Learning from Videos for 3D World: Enhancing MLLMs with 3D Vision Geometry Priors

### v2

@@ 5 Conclusion
We present a novel framework to enhance MLLMs’ 3D spatial understanding capability, which incorporates a 3D visual geometry encoder to provide latent 3D geometric information given only video inputs. While being straightforward, our extensive experiments show that our 4B model can outperform larger spatial-enhanced models on various 3D scene understanding tasks and spatial reasoning benchmarks, without relying on any explicit 3D scene input.
@@ Appendix A Experimental Setting and Details
A.1 Training Details.
Our model is built upon Qwen2.5-VL-3B [2], with a 3D geometry encoder initialized from VGGT-1B [43]. During training, the MLLM’s visual encoder, the integrated 3D geometry encoder, and the multimodal connector are frozen, while the MLLM backbone remains unfrozen with a learning rate of 5e-6. All experiments were conducted on 8 H100 80G GPUs. The training took 10 hours for 3D scene understanding and 12 hours for spatial reasoning instruction tuning.
A.2 Evaluation Details.
@@ Appendix B Data preparation
3D Visual Grounding ${{\bf X}}_{\texttt{User}:}$ Frame-0: <image>Frame-1:<image>Frame-2: <image>Frame-3: <image>$\cdots$
@@
[…]
@@ Appendix C Detailed Results on 3D Video Object Detection
Detailed quantitative results for our 3D video object detection are presented in Table 2. As illustrated in the table, our method significantly outperforms the baseline that does not utilize visual geometry across most categories. Nevertheless, we observe that detecting small objects such as pillows, lamps, and backpacks remains challenging.
4 Frames chair pillow cabinet table lamp couch desk stand bed backpack
Qwen2.5-VL-3B 37.7 11.9 10.2 35.0 4.4 39.0 33.7 23.3 64.8 27.4
+ Visual Geometry (VG LLM) 49.7 17.9 13.1 41.3 16.4 44.6 39.0 37.9 71.2 42.0
bathtub ottoman dresser bin toilet refrigerator stove microwave monitor computer
Qwen2.5-VL-3B 32.4 3.2 33.1 23.1 68.8 32.2 62.5 25.9 22.3 6.8
+ Visual Geometry (VG LLM) 33.5 3.8 40.3 39.2 83.4 57.3 68.8 32.1 29.1 3.8
6 Frames chair pillow cabinet table lamp couch desk stand bed backpack
Qwen2.5-VL-3B 32.8 11.4 7.8 31.3 5.6 32.2 31.6 20.6 58.8 28.9
+ Visual Geometry (VG LLM) 41.6 15.9 12.4 39.8 15.6 45.0 38.5 41.4 70.2 37.5
bathtub ottoman dresser bin toilet refrigerator stove microwave monitor computer
Qwen2.5-VL-3B 36.5 0 31.2 20.9 66.1 28.9 37.8 11.5 15.5 4.4
+ Visual Geometry (VG LLM) 33.8 2.8 41.1 33.1 80.6 64.3 68.5 21.1 22.1 3.1
Table 2: The detailed results on 3D video object detection. We report the F1 score for all categories.

### v3

@@ 5 Conclusion
We present a novel framework to enhance MLLMs’ 3D spatial understanding capability, which incorporates a 3D visual geometry encoder to provide latent 3D geometric information given only video inputs. While being straightforward, our extensive experiments show that our model can outperform larger spatial-enhanced models on various 3D scene understanding tasks and spatial reasoning benchmarks, without relying on any explicit 3D scene input.
Acknowledgements
This work is supported by National Key R&D Program of China (Project No. 2022ZD0161200, 2022ZD0161201). This work is also supported by Hong Kong Research Grant Council - Early Career Scheme (Grant No. 24200223).
@@ Appendix A Experimental Setting and Details
A.1 Evaluation Details.
@@ Appendix B Data preparation
[…]
@@
Text: There is a beige wooden bookshelf. placed next to another bookshelf. Output a JSON dictionary with the frame index in "frame" and its 3D bounding box in "box_3d" in the frame’s coordinates. ${{\bf X}}_{\texttt{Agent}:}$ ```json {"frame": 12, "bbox_3d": [-0.63, -0.83, 2.43, 3.0, 0.59, 2.35, -2.32, 1.18, 3.05]} ```   3D Dense Captioning ${{\bf X}}_{\texttt{User}:}$ $image><image>\cdots$ Carefully watch the video and describe the object located at [-0.89, -0.74, 2.45] in detail. ${{\bf X}}_{\texttt{Agent}:}$ A white cabinet in the corner of the room. in the direction from the door and from the inside. it will be on the left, there is a small brown table on the left side of the cabinet and a smaller table on the right side of the cabinet.   3D Video Object Detection ${{\bf X}}_{\texttt{User}:}$ $image><image>\cdots$ Detect the 3D bounding boxes in the camera coordinate system of the first frame. Output a json list where each entry contains the object name in "label" and its 3D bounding box in "box_3d". The 3D bounding box format should be [x_center, y_center, z_center, x_size, y_size, z_size, yaw, pitch, rolll]. ${{\bf X}}_{\texttt{Agent}:}$ ```json [    {"label": "bag", "bbox_3d": [0.0, -0.3, 1.0, 0.26, 0.26, 0.15, 1.67, 0.96, -2.98]}, $\qquad\cdots$ ] ```
@@ Appendix C Detailed Results on 3D Video Object Detection
Detailed quantitative results for our 3D video object detection are presented in Table 10 and 11. As illustrated in the table, our method significantly outperforms the baseline that does not utilize visual geometry across most categories. Nevertheless, we observe that detecting small objects such as pillows, lamps, and backpacks remains challenging.
Object
Qwen2.5-VL-3B
VG LLM-4B
Qwen2.5-VL-7B
VG LLM-8B
chair 32.8 41.6 36.1 48.7
pillow 11.4 15.9 12.4 18.9
cabinet 7.8 12.4 10.6 17.9
table 31.3 39.8 32.7 44.8
lamp 5.6 15.6 8.0 15.9
couch 32.2 45.0 40.7 46.4
desk 31.6 38.5 34.4 32.3
stand 20.6 41.4 31.3 47.3
bed 58.8 70.2 64.6 75.8
backpack 28.9 37.5 28.6 40.4
bathtub 36.5 33.8 38.4 45.2
ottoman 0.0 2.8 2.7 5.3
dresser 31.2 41.1 36.4 44.5
bin 20.9 33.1 25.0 38.5
toilet 66.1 80.6 68.6 83.2
refrigerator 28.9 64.3 33.2 62.0
stove 37.8 68.5 55.9 69.5
microwave 11.5 21.1 7.8 25.8
monitor 15.5 22.1 20.3 31.7
computer 4.4 3.1 5.2 11.6
Table 10: The detailed results on 3D video object detection in 6-frame setting. We report the F1 score for all categories.
Object
Qwen2.5-VL-3B
VG LLM-4B
Qwen2.5-VL-7B
VG LLM-8B
chair 37.7 49.7 41.2 54.0
pillow 11.9 17.9 13.2 22.9
cabinet 10.2 13.1 11.6 17.1
table 35.0 41.3 36.5 46.5
lamp 4.4 16.4 9.0 18.9
couch 39.0 44.6 41.1 47.0
desk 33.7 39.0 34.5 41.0
stand 23.3 37.9 33.7 46.1
bed 64.8 71.2 68.2 74.1
backpack 27.4 42.0 33.3 41.4
bathtub 32.4 33.5 36.6 42.1
ottoman 3.2 3.8 0.0 0.0
dresser 33.1 40.3 38.2 49.2
bin 23.1 39.2 30.2 39.8
toilet 68.8 83.4 68.7 82.5
refrigerator 32.2 57.3 38.5 53.6
stove 62.5 68.8 61.2 68.6
microwave 25.9 32.1 21.3 33.3
monitor 22.3 29.1 21.3 34.0
computer 6.8 3.8 12.0 11.4
Table 11: The detailed results on 3D video object detection in 4-frame setting. We report the F1 score for all categories.
@@ Appendix D More Visualization
Appendix E Limitations
While our model is built upon 3B and 7B MLLM backbones, its capabilities remain constrained by the model’s inherent capacity. Scaling up the model size could potentially enhance both fundamental abilities and generalization performance. Furthermore, this work focuses exclusively on supervised fine-tuning, whereas emerging research demonstrates that reinforcement learning can significantly enhance model reasoning capabilities. Since VGGT is trained on commonly used 3D scene datasets, evaluating it on benchmarks built from that same data (such as ScanNet) poses a greater risk of overestimating its performance compared to evaluating it on out-of-distribution benchmarks. We leave these for future work.
Appendix F Broader Impact
VG LLM is built upon Qwen2.5-VL and consequently shares many common challenges associated with multi-modal large language models (MLLMs). These challenges include hallucinations in visual understanding, inherited biases from the base models, and susceptibility to adversarial inputs. Despite these limitations, releasing VG LLM to the research community would be highly beneficial, as it could catalyze further advancements in 3D world understanding by leveraging Video LLMs and 3D vision geometry priors.